Find the capacitance of the Earth. The radius of the Earth is 6,400 km. By how much will the Earth's potential change if it receives a quantity of electricity equal to 1 C?



1
Expert's answer
2022-02-01T09:57:25-0500

The capacitance of shere

"C=4\pi\epsilon_0R\\
=4*3.14*8.85*10^{12}*6400*10^3=7.1*10^{-4}\:\rm F"

The Earth's potential

"V=\frac{q}{C}=\frac{7.1*10^{-4}\:\rm F}{1\:\rm C}=7.1*10^{-4}\:\rm V"